how do make x the subject in this equation:

y=25-e^x

Answer by stanbon(48569)

how do make x the subject in this equation:
y=25-e^x
----------------
solve for "x":
e^x = 25-y
x = ln(25-y)
==================
Cheers,
Stan H.